OpenStudy (anonymous):

solve by substitution: 2x+y=0 x-3y=-7 solve by elimination: 3x+4y=1 5x-3y=21

OpenStudy (anonymous):

for substitution solve for one variable i usually do y, 2x+y=0, subtract 2x over from both sides, y=-2x now we take -2x and plug it in for every y x-3(-2x)=-7 simplify to x+6x=-7 7x=-7 x=-1 and plug into the original y=-2x and y=2 that's substitution

OpenStudy (anonymous):

i was able to solve for y for the first time, but i wasnt able to solve for x

OpenStudy (anonymous):

when you solve for y you just put it into the second one for every you see, and solve for x, getting it in all x's makes it solvable

OpenStudy (anonymous):

i got down to x=-7+3y

OpenStudy (anonymous):

Let us solve the first system of equations by the method of substitution. 2x+y=0-----(1 x-3y=-7 i.e. x= 3y-7 ----(2 Now substituting x= 3y-7 in equation ---(1 we find 2(3y-7)+y=0 6y-14 +y=0 i.e. 7y-14 =0 i.e. 7y =14 i.e. y =14/7 i.e. y =2 Now substituting y =2 in any of the equation say ---(2 we find x= 3y-7 = 3*2-7= 6- 7 = -1 x=-1 Hence solution set is {-1, 2} @chevyguy

OpenStudy (anonymous):

you can do it that way to so now what you do is take that (-7-3y) and plug into the other for x, so you have 2(-7+3y) +y =0 does that part make sense?

OpenStudy (anonymous):

solving for x made sense thank you both!

OpenStudy (anonymous):

what about solving for elimination?

OpenStudy (anonymous):

Next, let us solve the second system of equations by the method of elimination. 3x+4y=1 -----(1 5x-3y=21 -----(2 Let us multiply eq (1 by 3 and eq (2 by 4 and adding (1 & (2 9x+12y=3 -----(1 20x-12y=84 -----(2 _________________________ 29 x +0 =87 i.e. 29 x =87 i.e. x =87/29 i.e. x =3 Now substitutingx =3 in any of the equation say ---(1 we find; 3*3+4y=1 i.e. 9+4y=1 i.e. 4y= 1-9 i.e. 4y= -8 i.e. y= -8/4 i.e. y= -2 hence solution set is {3, -2} @chevyguy

OpenStudy (anonymous):

@chevyguy Your problem has been solved.

OpenStudy (anonymous):

@dpasingh thank you so much! exactly what i needed was for someone to explain to me and show me what you did thank you so much!!!!!!!!!!

OpenStudy (anonymous):

@chevyguy Most welcome dear